When 120g of liquid L1 at 65°C was mixed with Yg of liquid L2 at 35°C the final temperature was 45°C. If the specific heat capacity of L2 is half that of L1 find the mass Yg
1
Expert's answer
2020-12-21T10:37:41-0500

Let's denote: "m_1 = 120g" is the mass of L1, "T_1 = 65\degree C" is the initial temperature of L1, "T_2 = 35\degree C" is the initial temperature of L2, "T_0 = 45\degree C" is the final temperature of the mixture, "c_2" - specific heat capacity of L2, "c_1 = 2c_2" - specific heat capacity of L1. "m_2 = Yg" is the mass of L2.

From the energy balance obtain:


"c_1m_1(T_1 - T_0) = c_2m_2(T_0-T_2)\\
m_2 = \dfrac{2c_2m_1(T_1 - T_0)}{c_2(T_0-T_2)}\\
m_2 = \dfrac{2m_1(T_1 - T_0)}{T_0-T_2}"

Numerically obtain:


"m_2 = \dfrac{2\cdot 120(65 - 45)}{45-35} = \dfrac{240\cdot 20}{10} = 480g"

Answer. 480g.